OpenStudy (anonymous):

What is the product of the following matrices? [ 8 -3] x [ 4 7 ] -1 2 6 0

myininaya (myininaya):

\[\left[\begin{matrix}8 & -3 \\ 6 & 0\end{matrix}\right]*\left[\begin{matrix}4 & 7 \\ -1 & 2\end{matrix}\right]\] ?

myininaya (myininaya):

it looks like you have a 1 by 2 times 3 by 2 but i don't think you mean that since you can't multiply a 1*2 times a 3 by 2

OpenStudy (anonymous):

so its undefined?

myininaya (myininaya):

if you have a 1 by 2 times a 3 by 2 yes

myininaya (myininaya):

\[\left[\begin{matrix}8 & -3 \end{matrix}\right]*\left[\begin{matrix}4 & 7 \\ -1 & 2\\6 & 0\end{matrix}\right]\] is this what you have?

myininaya (myininaya):

you can't multiply these if so

OpenStudy (anonymous):

yes. It's not possible, therefore the product is undefined?